# Unsupervised Morphological Tree Tokenizer

### v1

@@ 5 Conclusion
In this work, we proposed a tree-based tokenizer (TreeTok) to better align with morphology. To make the method as general-purpose as possible, the most critical challenge is how to induce character-level structures without relying on human-annotated data. Our research revealed that the key to solving this problem lies in recognizing the indecomposability of morphemes. We introduce a composition model with a MorphOverriding mechanism to incorporate this inductive bias and propose two self-supervised objectives. Together, we can effectively induce tree structures that are highly consistent with human-labeled morphology. Ultimately, we achieved consistently positive results when comparing TreeTok against strong baselines like BPE and WordPiece on various tasks. This study provides new insights into unsupervised morphological segmentation and offers a novel approach different from traditional statistical models.
@@ 6 Limitations
Our main limitation is that we need additional training and inference overheads. Considering that the composition model only needs to be trained once and the overall time consumption is acceptable5551 day for 8 A100 for WikiText-103, we believe it is not a fatal flaw. Regarding inference cost, because a lightweight parser is produced as a byproduct, it can be afforded by even CPU environments.

### v2

@@ 5 Conclusion
Our work introduces the first effective unsupervised neural model for character-level structure induction. We discovered that recognizing the indecomposability of morphemes is key, and to address this, we developed a composition model with a MorphOverriding mechanism alongside two self-supervised objectives. TreeTok induces tree structures that closely match human-labeled morphology and consistently outperforms baselines like BPE and WordPiece across various tasks, offering new insights into unsupervised morphological segmentation.
@@ 6 Limitations
| Tokenizer | Avg Time/Token (s) |
| BPE | 2.49e-05 |
| WordPiece | 2.33e-05 |
| Unigram | 2.54e-05 |
| TreeTok (single processing) | 1.98e-03 |
Our main limitation is that we need additional training and inference overheads. Considering that the composition model only needs to be trained once and the overall time consumption is acceptable666less than 1 day for 8 A100 for WikiText-103, we believe it is not a fatal flaw. Regarding inference cost, because a lightweight parser is produced as a byproduct, it can be afforded by even CPU environments. According to Table 5, Treetok’s average processing time per token is longer than other tokenizers. However, if we allow Treetok to tokenize in batches on a GPU in advance, this gap can be easily compensated. Furthermore, we can maintain a cache of high-frequency words to avoid repeated tokenization. In wikitext-103, the hit rate for a cache that stores the top 100000 frequent words is 98.11%, which means only 2% tokens need to be parsed on the fly. e.g. for 1000 tokens, it only needs 0.16s to parse. Furthermore, these results are based on single-core computation, and there is still room for multi-core acceleration.
7 Acknowledgements
QZ acknowledges support through the NSF under award 1922658.